A 4.0 kg box is on a frictionless 35° slope and is connected via a massless string over a massless, frictionless pulley to a hanging 2.0 kg weight. What is the acceleration (magnitude and direction) of the 2.0 kg block as the blocks are moving?
Answer #1

FBD of 2 kg

m2 a = m2 g - T

T = m2 g - m2 a... (i)

FBD of mass placed at slope

m1 a = T - m1 g sin x

puttinh value of T in above equation

m1 a = m2 g - m2 a - m1 g sin x

a ( m1 + m2) = m2 g - m1 g sin x

a ( 4 + 2) = 2* 9.8 - 4* 9.8* sin 35

a = 0.4807 m/s^2

directed upwards (towards the pulley)

=======

Comment before rate in case any doubt, will reply for sure.. goodluck